What mass of coppef would you be able to raise from 92 C to 192 C py applying 4.100 joules of heal?


1
Expert's answer
2022-04-28T15:16:03-0400

Solution:

q = m × C × ΔT

where:

q = amount of heat energy gained or lost by substance (J)

m = mass of sample (g)

C = specific heat capacity (J °C−1 g−1)

ΔT = change in temperature (°C)


q = 4100 J

m = ??? 

C = 0.385 J °C−1 g−1

ΔT = 192°C − 92°C = 100°C


Thus:

m = q / (C × ΔT)

m = (4100 J) / (0.385 J °C−1 g−1 × 100°C) = 106.4935 g = 106.5 g

Mass of copper = 106.5 g


Answer: 106.5 grams of copper